OpenStudy (anonymous):

Form a polynomial, f(x), with real coefficients having the given degree and zeros. Degree: 4; zeros: 3+2i; 4, multiplicity 2

OpenStudy (aum):

\[ x = 3 + 2i \\ x - 3 = 2i \\ (x-3)^2 = (2i)^2 = -4 \\ x^2 - 6x + 9 = - 4 \\ x^2 - 6x + 13 = 0 \\ \text{ } \\ x = 4 \\ (x-4) = 0 \\ \text{multiplicity of 2 makes it:} (x-4)^2 \\ f(x) = (x^2 - 6x + 13)(x-4)^2 \\ \text{Multiply it out and arrange it in decreasing powers of x} \]